A 0.50 kg ball traveling at 6.0 m/s collides head-on with a 1.00-kg ball moving in the
opposite direction at a velocity of –12.0 m/s.
The 0.50 kg ball moves away at –14 m/s after
the collision. Find the velocity of the second ball.
1
Expert's answer
2011-08-25T13:19:41-0400
Let m1 be the mass of the first ball (m1 = 0.5 kg), m2 - the mass of the second ball (m2 = 1 kg).
U1, V1 - speeds of the first ball before and after the collision, U2, V2 - speeds of the second ball before and after the collision.
According to the law of conservation of linear momentum (let the initial direction of the first ball be positive):


m1U1 - m2U2=-m1V1 + m2V2


V2 = (U1 + V1)*m1/m2 - U2
V2 = (6 + 14)*0.5/1 - 12 = -2 m/s.


So, the second ball continued moving at the same direction at 2 m/s.